# An Analysis of Concept Bottleneck Models: Measuring, Understanding, and Mitigating the Impact of Noisy Annotations

### v1

@@ D.5 Evaluating Concept Noise in LLM-Generated Annotations
As shown in Table 8, the resulting annotations agree with human-labeled ground truth only 21.8% of the time, revealing a significant misalignment. This highlights a key issue, i.e., general-purpose LLMs often produce annotations that diverge from expert standards, introducing a new source of noise. To assess the effect of this noise, we trained a CBM on the LLM-annotated dataset. This model achieved 68.5% accuracy, a 5.8% drop from the baseline trained on expert annotations. Unlike traditional concept noise, which often stems from human inconsistency, LLM-induced noise may arise from the the limited domain grounding and generalization of model. These findings expose the current limitations of LLM-generated concept supervision and reinforce the broader challenge of label noise in building robust and interpretable CBMs, particularly in high-stakes domains.
D.6 Additional Evidence of Disrupted Representation Alignment
(a) Brewer Blackbird
(b) Importance shift in key concepts
(c) Key concepts under 40% noise
In Section 3.2, we illustrated how noise disrupts the representation alignment for the class Le Conte Sparrow. Here, we provide additional qualitative results for the class Brewer Blackbird, which demonstrates a comparable degradation pattern.
As shown in Figure 15, at a moderate noise level (i.e., 20%), the relative ordering of predictive dimensions remains quite well preserved. In contrast, at higher noise levels (i.e., 40%), the saliency of informative dimensions deteriorates significantly, while irrelevant or spurious concepts become disproportionately influential. These findings further substantiate our conclusion that noise induces representational misalignment, and when this misalignment coincides with susceptible concepts, it becomes a key driver of instability in CBMs under noisy supervision.
@@ D.7 Effect of Target Predictor Complexity
As shown in Table 9, the non-linear variant consistently underperforms its linear counterpart, with performance degrading more severely as noise increases. These results suggest that non-linear predictors are more susceptible to overfitting noisy concept representations, thereby undermining generalization. Due to the concept bottleneck imposed by CBMs, increasing the complexity of $f$ does not inherently enhance predictive accuracy; rather, it can exacerbate the propagation of noise from the concept layer to the output. Overall, these findings indicate that under noisy conditions, simpler target predictors may offer greater robustness. However, this conclusion is drawn from a preliminary comparison between a linear and a two-layer model. Exploring a broader range of non-linear architectures and incorporating regularization techniques may offer valuable insights into the trade-off between expressiveness and robustness.
@@ D.8 Impact of Noise on Intervention Effectiveness
As outlined in Section C.5, we evaluate six concept selection strategies for intervention: Random, UCP, CCTP, LCP, ECTP, and EUDTP. While the main paper focuses on the first three, here we provide an extended analysis of the remaining strategies (e.g., LCP, ECTP, and EUDTP) for a more comprehensive evaluation.
@@
Figure 16 presents task accuracy improvements achieved through concept-level interventions on the CUB and AwA2 datasets under varying levels of noise. Across both datasets, we observe a consistent pattern: the effectiveness of interventions diminishes as the noise level increases, irrespective of the intervention selection strategy. These findings corroborate our main results and further demonstrate that noise significantly compromises the reliability of concept-based interventions. Moreover, for all intervention methods, the upper bound on task accuracy (which corresponds to interventions applied to all concepts) systematically decreases with increasing noise levels. This downward trend underscores the importance of incorporating robustness-aware selection criteria when designing intervention strategies for noisy environments.
@@ D.9 Comparison of Prediction Accuracy between SAM and Base Optimizer
Concept Acc 93.8$\pm$$0.1$ 91.6$\pm$$0.0$ 89.1$\pm$$0.0$ 85.4$\pm$$0.0$
Base Task Acc 57.7$\pm$$2.0$ 50.3$\pm$$0.7$ 23.3$\pm$$1.2$ 4.0$\pm$$0.7$
Concept Acc 94.6$\pm$$0.1$ 92.5$\pm$$0.1$ 89.7$\pm$$0.1$ 86.3$\pm$$0.1$ +0.8
Independent SAM Task Acc 61.8$\pm$$1.8$ 54.2$\pm$$0.7$ 28.5$\pm$$1.4$ 5.0$\pm$$1.4$ +3.6
Concept Acc 93.8$\pm$$0.0$ 91.6$\pm$$0.0$ 89.1$\pm$$0.0$ 85.4$\pm$$0.1$
Base Task Acc 66.6$\pm$$0.4$ 59.3$\pm$$0.6$ 47.0$\pm$$1.7$ 6.1$\pm$$2.6$
Concept Acc 94.6$\pm$$0.1$ 92.5$\pm$$0.1$ 89.7$\pm$$0.1$ 86.3$\pm$$0.1$ +0.8
Sequential SAM Task Acc 70.5$\pm$$0.6$ 63.5$\pm$$0.9$ 50.1$\pm$$1.1$ 10.7$\pm$$6.0$ +4.0
Concept Acc 85.9$\pm$$0.5$ 78.4$\pm$$0.6$ 67.6$\pm$$1.2$ 57.3$\pm$$0.3$
Base Task Acc 75.2$\pm$$0.3$ 69.2$\pm$$0.5$ 59.8$\pm$$0.3$ 50.1$\pm$$0.5$
Concept Acc 86.0$\pm$$0.2$ 78.5$\pm$$0.1$ 68.0$\pm$$0.8$ 57.9$\pm$$0.3$ +0.3
Joint SAM Task Acc 76.1$\pm$$0.4$ 69.9$\pm$$0.6$ 60.8$\pm$$0.4$ 50.6$\pm$$1.5$ +0.8
@@
Concept Acc 78.4$\pm$$0.8$ 78.1$\pm$$0.7$ 77.3$\pm$$0.3$ 75.3$\pm$$0.8$
Base Task Acc 85.5$\pm$$0.3$ 82.3$\pm$$1.4$ 77.1$\pm$$0.5$ 41.9$\pm$$1.0$
Concept Acc 78.6$\pm$$0.7$ 78.5$\pm$$0.5$ 77.9$\pm$$0.7$ 75.9$\pm$$1.3$ +0.5
Independent SAM Task Acc 88.1$\pm$$0.1$ 85.7$\pm$$0.4$ 78.6$\pm$$2.8$ 46.5$\pm$$1.4$ +3.0
Concept Acc 78.4$\pm$$0.8$ 78.1$\pm$$0.7$ 77.3$\pm$$0.3$ 75.3$\pm$$0.8$
Base Task Acc 87.6$\pm$$0.3$ 85.8$\pm$$0.3$ 81.8$\pm$$1.1$ 70.1$\pm$$3.9$
Concept Acc 78.6$\pm$$0.7$ 78.5$\pm$$0.5$ 77.9$\pm$$0.7$ 75.9$\pm$$1.3$ +0.5
Sequential SAM Task Acc 89.5$\pm$$0.5$ 88.0$\pm$$0.5$ 82.6$\pm$$3.1$ 69.6$\pm$$6.3$ +1.1
Concept Acc 74.2$\pm$$0.4$ 70.1$\pm$$0.8$ 65.4$\pm$$0.3$ 57.4$\pm$$0.2$
Base Task Acc 84.2$\pm$$0.1$ 83.0$\pm$$0.3$ 82.2$\pm$$0.1$ 81.7$\pm$$0.3$
Concept Acc 74.9$\pm$$0.4$ 72.7$\pm$$0.6$ 67.7$\pm$$0.7$ 58.9$\pm$$0.9$ +1.8
Joint SAM Task Acc 89.0$\pm$$0.1$ 88.4$\pm$$0.2$ 87.3$\pm$$0.2$ 86.6$\pm$$0.3$ +5.1
Table 10 reports the overall task and concept prediction accuracy of CBMs trained with SAM and the baseline (Base) on the CUB and AwA2 datasets under various noise settings ($\gamma\in\{0.0,0.2,0.4\}$). Across all noise levels and types, SAM consistently outperforms Base in both concept and task accuracy, demonstrating its robustness to noise.
@@ D.10 Robustness of CBMs across Backbone Architectures
Table 11 presents the task and concept prediction accuracies for each configuration. Our findings indicate that, regardless of the backbone architecture, CBMs experience significant performance degradation as noise increases. However, models trained with the SAM optimizer consistently exhibit improved robustness compared to those trained with SGD. Specifically, SAM yields average task accuracy improvements of 5.07% for ResNet-18 and 7.31% for ViT-B/16 under noisy conditions.
@@
D.11 Concept-Level Qualitative Results
In Section 2.3, we presented qualitative t-SNE [49] visualizations to illustrate how noise deteriorates the interpretability of learned concept representations. Here, we provide additional qualitative results focusing on a subset of representative concepts related to bird morphology and coloration. Specifically, we examine the spatial clustering patterns of concept predictions for various bill shapes (e.g., Curved Bill, Spatulate Bill, and All Purpose Bill) as well as wing colors (e.g., Blue Wings and Orange Wings). These concepts were selected based on their semantic diversity and prevalence in the dataset. The resulting visualizations further confirm that increased noise disrupts the cluster structure of concept embeddings, reducing the ability of model to maintain separable and interpretable representations.

### v2

@@ D.6 Evaluating Concept Noise in LLM-Generated Annotations
As shown in Table˜8, the resulting annotations agree with human-labeled ground truth only 21.8% of the time, revealing a significant misalignment. This highlights a key issue, i.e., general-purpose LLMs often produce annotations that diverge from expert standards, introducing a new source of noise. To assess the effect of this noise, we trained a CBM on the LLM-annotated dataset. This model achieved 68.5% accuracy, a 5.8% drop from the baseline trained on expert annotations. Unlike traditional concept noise, which often stems from human inconsistency, LLM-induced noise may arise from the the limited domain grounding and generalization of model. These findings expose the current limitations of LLM-generated concept supervision and reinforce the broader challenge of label noise in building robust and interpretable CBMs, particularly in high-stakes domains.
@@ D.7 Effect of Target Predictor Complexity
As shown in Table˜9, the non-linear variant consistently underperforms its linear counterpart, with performance degrading more severely as noise increases. These results suggest that non-linear predictors are more susceptible to overfitting noisy concept representations, thereby undermining generalization. Due to the concept bottleneck imposed by CBMs, increasing the complexity of $f$ does not inherently enhance predictive accuracy; rather, it can exacerbate the propagation of noise from the concept layer to the output. Overall, these findings indicate that under noisy conditions, simpler target predictors may offer greater robustness. However, this conclusion is drawn from a preliminary comparison between a linear and a two-layer model. Exploring a broader range of non-linear architectures and incorporating regularization techniques may offer valuable insights into the trade-off between expressiveness and robustness.
@@ D.8 Impact of Noise on Intervention Effectiveness
As outlined in Section˜C.5, we evaluate six concept selection strategies for intervention: Random, UCP, CCTP, LCP, ECTP, and EUDTP. While the main paper focuses on the first three, here we provide an extended analysis of the remaining strategies (e.g., LCP, ECTP, and EUDTP) for a more comprehensive evaluation.
@@
Figure˜17 presents task accuracy improvements achieved through concept-level interventions on the CUB and AwA2 datasets under varying levels of noise. Across both datasets, we observe a consistent pattern: the effectiveness of interventions diminishes as the noise level increases, irrespective of the intervention selection strategy. These findings corroborate our main results and further demonstrate that noise significantly compromises the reliability of concept-based interventions. Moreover, for all intervention methods, the upper bound on task accuracy (which corresponds to interventions applied to all concepts) systematically decreases with increasing noise levels. This downward trend underscores the importance of incorporating robustness-aware selection criteria when designing intervention strategies for noisy environments.
@@ D.9 Comparison of Prediction Accuracy between SAM and Base Optimizer
[…]
@@
[…]
Table˜10 reports the overall task and concept prediction accuracy of CBMs trained with SAM and the baseline (Base) on the CUB and AwA2 datasets under various noise settings ($\gamma\in\{0.0,0.2,0.4\}$). Across all noise levels and types, SAM consistently outperforms Base in both concept and task accuracy, demonstrating its robustness to noise.
@@ D.10 Robustness of CBMs across Backbone Architectures
Table˜11 presents the task and concept prediction accuracies for each configuration. Our findings indicate that, regardless of the backbone architecture, CBMs experience significant performance degradation as noise increases. However, models trained with the SAM optimizer consistently exhibit improved robustness compared to those trained with SGD. Specifically, SAM yields average task accuracy improvements of 5.07% for ResNet-18 and 7.31% for ViT-B/16 under noisy conditions.
@@
D.11 Additional Evidence that Susceptibility Approximates Uncertainty
In Section˜4.2, we demonstrated on the CUB dataset that susceptibility can be effectively approximated by uncertainty. Here, we provide additional results on the AwA2 dataset, which exhibit similar trends.
As shown in Figure˜18, a positive correlation between uncertainty and susceptibility consistently appears across datasets. While the AwA2 dataset shows a mild negative correlation within the non-susceptible set, this does not contradict the strong positive relationship observed among susceptible concepts. These findings support that our analysis captures a general and dataset-independent pattern.
(a) $\gamma=0.1$
(b) $\gamma=0.2$
(c) $\gamma=0.3$
Figure 18: Correlation between uncertainty and susceptibility on AwA2. As with the CUB dataset, Pearson correlation coefficients ($\rho$) are reported separately for susceptible and non-susceptible sets, showing a positive correlation in the former. A negative correlation appears in the non-susceptible subset but does not alter the positive relationship observed for the susceptible subset.
D.12 Concept-Level Qualitative Results
In Section˜2.3, we presented qualitative t-SNE [51] visualizations to illustrate how noise deteriorates the interpretability of learned concept representations. Here, we provide additional qualitative results focusing on a subset of representative concepts related to bird morphology and coloration. Specifically, we examine the spatial clustering patterns of concept predictions for various bill shapes (e.g., Curved Bill, Spatulate Bill, and All Purpose Bill) as well as wing colors (e.g., Blue Wings and Orange Wings). These concepts were selected based on their semantic diversity and prevalence in the dataset. The resulting visualizations further confirm that increased noise disrupts the cluster structure of concept embeddings, reducing the ability of model to maintain separable and interpretable representations.
@@ D.12 Concept-Level Qualitative Results
D.13 Noise Injection Protocol
In Figure˜6, noise is injected at the concept level with different random seeds, which can substantially influence how the results are interpreted. To assess whether the same concepts are consistently identified as susceptible across different noise realizations (or whether susceptibility varies randomly) we report the susceptible concepts obtained for each seed in Figure˜20. We find that similar concepts are repeatedly identified as susceptible across seeds, indicating structural sensitivity rather than random variation.
Figure 20: Susceptible concept sets across random seeds. Susceptible concept sets are identified for each random seed, and similar concepts are consistently detected across seeds.